An airplane keeps a speed of 635 km/h relative to the air it is flying through, as it makes a trip to a city 

765 km away to the north. (a) What time interval is required for the trip if the plane flies through a 

headwind blowing at 37.5 km/h toward the south? (b) What time interval is required if there is a tailwind 

with the same speed? (c) What time interval is required if there is a crosswind blowing at 37.5 km/h to 

the east relative to the ground? 


1
Expert's answer
2021-02-15T17:42:29-0500

a)


"t=\frac{765}{635-37.5}=1.28\ h"

b)


"t=\frac{765}{635+37.5}=1.14\ h"

c)


"\theta=\arctan{\frac{37.5}{635}}=3.38\degree"

"t=\frac{765}{635\cos{3.38}}=1.21\ h"